I need help on this problem 6v-9/3=v.
Answer by checkley77(12844)

6v-9/3=v
6v-v=9/3
5v=3
v=3/5 ans.
Proof:
6*3/5-3=3/5
18/5-3=3/5
18/5-15/5=3/5
3/5=3/5